please help

the length of a rectangular playing field is 5 ft less than twice its width. If the perimeter of the playing field is 230 ft. find the length and width of the field.

Answer by checkley71(8403)

L=2W-5
2L+2W=230
2(2W-5)+2W=230
4W-10+2W=230
6W=230+10
6W=240
W=240/6
W=40 ANSWER FOR THE WIDTH.
L=2*40-5
L=80-5
L=75 ANSWER FOR THE LENGTH.
PROOF
2*75+2*40=230
150+80=230
230=230